2.7 Find the power dissipated in the following resistors

  1. 100kΩ resistor with a 10A of current
  2. 45Ω resistor with a 25mA of current
  3. 1kΩ resistor with a 75mA of current
1
Expert's answer
2021-03-17T16:54:51-0400

1) "P_1=I_1^2R_1=10^2\cdot100000=10\cdot10^6\ (J)=10\ MJ"


2) "P_2=I_2^2R_2=0.025^2\cdot45=0.028\ (J)"


3) "P_3=I_3^2R_3=0.075^2\cdot1000=5.625\ (J)"